OpenStudy (anonymous):

1,10,19,28... help medals wbg

OpenStudy (anonymous):

addition of 9

OpenStudy (anonymous):

It appears to be an arithmetic sequence, given by a(n)=9n-8, if we take a(1) as one.

OpenStudy (anonymous):

1+9=10+9=19+9=28....